OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = Cube root of quantity x minus four.

OpenStudy (anonymous):

you want to replace the f(x) with y which makes it simpler to solve. Then you want to switch the two variables (x and y) . After you do that you want to get y by itself. once you have done that, remember that our y variable stands for f(x)

OpenStudy (anonymous):

i think it wants you to compose the functions \[f(x)=x^3+4,g(x)=\sqrt[3]{x-4}\] so \[f(g(x))=\sqrt[3]{x^3+4-4}\] is a start when you get \(f(g(x))=x\) you are done

OpenStudy (anonymous):

oh i missed that part. your right